OpenStudy (anonymous):

Write a polynomial of the smallest degree with roots 1, -3, and 4.

OpenStudy (anonymous):

just write it out in factored form, then multiply them together one by one (x-4)(x+2)(x-1)(x-2) start expanding

OpenStudy (anonymous):

f(x) = (x-4)(x+2)(x-1)(x-2)

OpenStudy (leader):

f(x)=(x-4)(x+3)(x-1) f(x)=\(\ x^3-2x^2-11x+12\)

OpenStudy (leader):

@josue12372 @timanti